Age of father and son is 3:10. Father is 56 years older than his son find their ages
Found 2 solutions by stanbon, Alan3354:
Answer by stanbon(75887)

Age of father and son is 3:10. Father is 56 years older than his son find their ages
------
f = s + 56
3f = 10s
---------
(10/3)s + s = 56
10s = 3s = 168
--
13s = 168
s = 12.92 = 12 yrs 11 mts
father = (10/3)12.92 = 43.067 = 43 yrs 1 mth
--------------
Cheers,
Stan H.

Answer by Alan3354(69443)

Age of father and son is 3:10. Father is 56 years older than his son find their ages
---------------
Their ages are 3x and 10x
10x - 3x = 7x = 56
x = 8
-----
--> 24 and 80